# Improved Particle Approximation Error for Mean Field Neural Networks

### v1

@@ Abstract
Mean-field Langevin dynamics (MFLD) minimizes an entropy-regularized nonlinear convex functional defined over the space of probability distributions. MFLD has gained attention due to its connection with noisy gradient descent for mean-field two-layer neural networks. Unlike standard Langevin dynamics, the nonlinearity of the objective functional induces particle interactions, necessitating multiple particles to approximate the dynamics in a finite-particle setting. Recent works (Chen et al.,, 2022; Suzuki et al., 2023b, ) have demonstrated the uniform-in-time propagation of chaos for MFLD, showing that the gap between the particle system and its mean-field limit uniformly shrinks over time as the number of particles increases. In this work, we improve the dependence on logarithmic Sobolev inequality (LSI) constants in their particle approximation errors, which can exponentially deteriorate with the regularization coefficient. Specifically, we establish an LSI-constant-free particle approximation error concerning the objective gap by leveraging the problem structure in risk minimization. As the application, we demonstrate improved convergence of MFLD, sampling guarantee for the mean-field stationary distribution, and uniform-in-time Wasserstein propagation of chaos in terms of particle complexity.
@@ 3.2.3 Uniform-in-time Wasserstein propagation of chaos
Proof techniques and key results
In this section, we provide the proof sketch of Theorem 1. Our analysis carefully treats the non-linearity of $F_{0}$ because the differences between $\mathcal{L}$ and $\mathcal{L}^{(N)}$, and between their solutions $\mu_{*}$ and $\mu^{(N)}_{*}$ come from this non-linearity. In fact if $F_{0}$ is a linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]$ ($\exists f:\mathbb{R}^{d}\rightarrow\mathbb{R}$), then $\mathcal{L}^{(N)}$ becomes simply some of $N$-independent objectives: $\mathcal{L}^{(N)}(\mu^{(N)})=\mathcal{L}^{(N)}(\mu^{(N)}_{1},\ldots,\mu^{(N)}_{N})=\sum_{i=1}^{N}\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i}}[f(X^{i})+\lambda^{\prime}\|X^{i}\|_{2}^{2}]+\lambda\sum_{i=1}^{N}\mathrm{Ent}(\mu^{(N)}_{i})=\sum_{i=1}^{N}\mathcal{L}(\mu^{(N)}_{i})$, where $\mu^{(N)}_{i}$ are marginal distributions on $X^{i}$. This results in $\mathcal{L}^{(N)}(\mu^{(N)}_{*})=N\mathcal{L}(\mu_{*})$ and $\mu^{(N)}_{*}=\mu^{\otimes N}_{*}$, and thus there is no approximation error by using finite-particles as also deduced from Theorem 1 with $L=0$. Therefore, we should take into account the non-linearity of $F_{0}$ to tightly evaluate the gap between $\mathcal{L}$ and $\mathcal{L}^{(N)}$.
@@
Intuitively, $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ is small because the empirical distribution $\mu_{\mathbf{X}}$ ($\mathbf{X}\sim\mu^{\otimes N}_{*}$) converges to $\mu_{*}$ by law of large numbers. In fact, we can reduce the evaluation of this term to the application of a concentration inequality, yielding a bound: $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq\frac{2LR^{2}}{N}$ (See the proof of Theorem 1) in Appendix 5. Then, we arrive at Theorem 1.
Conclusion and Disuccion
@@ 5.1 Finite-particle approximation error
As discussed in Section 4, the evaluation of $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ completes the proof. For any function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we have

### v2

@@ Abstract
Mean-field Langevin dynamics (MFLD) minimizes an entropy-regularized nonlinear convex functional defined over the space of probability distributions. MFLD has gained attention due to its connection with noisy gradient descent for mean-field two-layer neural networks. Unlike standard Langevin dynamics, the nonlinearity of the objective functional induces particle interactions, necessitating multiple particles to approximate the dynamics in a finite-particle setting. Recent works (Chen et al.,, 2022; Suzuki et al., 2023b, ) have demonstrated the uniform-in-time propagation of chaos for MFLD, showing that the gap between the particle system and its mean-field limit uniformly shrinks over time as the number of particles increases. In this work, we improve the dependence on logarithmic Sobolev inequality (LSI) constants in their particle approximation errors which can exponentially deteriorate with the regularization coefficient. Specifically, we establish an LSI-constant-free particle approximation error concerning the objective gap by leveraging the problem structure in risk minimization. As the application, we demonstrate improved convergence of MFLD, sampling guarantee for the mean-field stationary distribution, and uniform-in-time Wasserstein propagation of chaos in terms of particle complexity.
@@ 3.2.3 Uniform-in-time Wasserstein propagation of chaos
Proof outline and key results
In this section, we provide the proof sketch of Theorem 1. Our analysis carefully treats the non-linearity of $F_{0}$ because the differences between $\mathcal{L}$ and $\mathcal{L}^{(N)}$, and between their solutions $\mu_{*}$ and $\mu^{(N)}_{*}$ come from this non-linearity. In fact if $F_{0}$ is a linear functional: $F_{0}(\mu)=\mathbb{E}_{\mu}[f]$ ($\exists f:\mathbb{R}^{d}\rightarrow\mathbb{R}$), then $\mathcal{L}^{(N)}(\mu^{(N)})=\sum_{i=1}^{N}\mathbb{E}_{X^{i}\sim\mu^{(N)}_{i}}[f(X^{i})+\lambda^{\prime}\|X^{i}\|_{2}^{2}]+\lambda\mathrm{Ent}(\mu^{(N)})\geq\sum_{i=1}^{N}\mathcal{L}(\mu_{i}^{(N)})$, where $\mu^{(N)}_{i}$ are marginal distributions on $X^{i}$. This results in $\mu^{(N)}_{*}=\mu^{\otimes N}_{*}$ and $\mathcal{L}^{(N)}(\mu^{(N)}_{*})=N\mathcal{L}(\mu_{*})$, and thus there is no approximation error by using finite-particles as also deduced from Theorem 1 with $L=0$. Therefore, we should take into account the non-linearity of $F_{0}$ to tightly evaluate the gap between $\mathcal{L}$ and $\mathcal{L}^{(N)}$.
@@
Intuitively, $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ is small because the empirical distribution $\mu_{\mathbf{X}}$ ($\mathbf{X}\sim\mu^{\otimes N}_{*}$) converges to $\mu_{*}$ by law of large numbers. In fact, we can reduce the evaluation of this term to the application of a concentration inequality, yielding a bound: $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]\leq\frac{2LR^{2}}{N}$ (see the proof of Theorem 1 in Appendix 5). Then, we arrive at Theorem 1.
Conclusion and Discussion
@@ 5.1 Finite-particle approximation error
As discussed in Section 4, the evaluation of $\mathbb{E}_{\mathbf{X}\sim\mu^{\otimes N}_{*}}\left[B_{F}(\mu_{\mathbf{X}},\mu_{*})\right]$ completes the proof. For any function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ so that the following integral is well defined, we have